OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 3t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (ivancsc1996):

You need to differentiate the equation.

OpenStudy (ivancsc1996):

\[s'(t)=-3\]

OpenStudy (jhannybean):

instantaneous velocity is the function of acceleration. The second derivative. \[\large \frac{d}{dx} [s(t)]\]

OpenStudy (jhannybean):

what is the derivative of this?

OpenStudy (anonymous):

d/dx(-3)

OpenStudy (ivancsc1996):

No, intantaneus velocity is the first derivativo of position.

OpenStudy (ivancsc1996):

No, the answer is -3.

OpenStudy (anonymous):

wait the differentiate thing... i need help with another type like this

OpenStudy (jhannybean):

hhh you're right! ok.

OpenStudy (anonymous):

a few actually -.-

OpenStudy (anonymous):

Find the derivative of f(x) = -10x2 + 4x at x = 11

OpenStudy (anonymous):

-216 -196 -176 -363

OpenStudy (ivancsc1996):

I think so. Velocity is the rate of change of position, therefore, the instantaneous velocity is just the derivative of the position function. Since s(x) is a linear function the the derivative s'(x) is a constant (-3). So the answer is -3. That meanse that the velocity is always -3.

OpenStudy (jhannybean):

d/dx (-10x^2) + d/dx(4x then plug in x.

OpenStudy (anonymous):

-10(11)^2 + 4(11)

OpenStudy (jhannybean):

yes you were right,instantaneous velocity is approximately equal to the rate of change of an object at a point.

OpenStudy (jhannybean):

no. no.

OpenStudy (jhannybean):

\[\large f'(x) = -10(2)x +4\]

OpenStudy (anonymous):

-20(11) + 4

OpenStudy (anonymous):

-216, got it! Thanks jhanny and ivan

OpenStudy (jhannybean):

Np.

OpenStudy (anonymous):

Oh, I've got 2 more -.-

OpenStudy (anonymous):

can I post a new Q and tag you two?

OpenStudy (ivancsc1996):

\[\frac{ d }{ dx }(-10x ^{2}+4x)=-10\frac{ d }{ dx }x ^{2}+4\frac{ d }{ dx }x=-20x+4\]

OpenStudy (jhannybean):

haha sure.